The legs of a right triangle have lengths A and B satisfying A + B = 10. Which values of A and B maximize the area of the triangle?
1
Expert's answer
2019-10-27T06:49:03-0400

Solution:

We are going to find the maximum value of Area of Triangle.


Given,

The legs of a right triangle have lengths A and B


"A + B = 10"





Here Base = A and Height = B

We have a formula for area of the triangle, that is


Area of the Right triangle ="\frac {1} {2} \times Base \times height"

= "\frac {1} {2} \times A \times B"

Now we can solvve for A and B using the concept A.M (Arithmetic mean) and G.M (Geometric mean)


"A.M \space of \space A \space and \space B = \frac {A+B} {2}"


"G.M \space of \space A \space and \space B = \sqrt {AB}"


We know,

"G.M \le A.M"

"\sqrt {AB}\le \frac {A+B} {2}"


"\sqrt {AB}\le \frac {10} {2} \\\sqrt {AB}\le 5"

"AB \le 25"

Maximum Area would be = "\frac {1}{2} AB = 12.5", if the A and B are equal

So, A = B = 5.


Answer: A = B = 5